OpenStudy (anonymous):

derivative of (3+cos^3(3x))^-1/3

OpenStudy (anonymous):

-1/3(3+cos^3(3x))^-4/3, this is the first step in the chain rule for this problem. are you familiar with the chain rule?

OpenStudy (anonymous):

\[\frac{ -1 }{ 3 } (\cos ^{3}(3x))^{-4/3}(-3\sin ^{2}(3x))(3)\] This is the complete chain rule for this problem, remember that if F=f(gx) then F'=(f'(gx))(g'(x))

OpenStudy (anonymous):

but my professor gave me this as the answer.